OpenStudy (anonymous):

What is the lim as x approaches zero sinx/x^2+3x?

zepdrix (zepdrix):

Is that all in the denominator? (x^2+3x)

OpenStudy (anonymous):

Yes that is all.

zepdrix (zepdrix):

Factor the denominator: \[\Large\bf\sf \lim_{x\to0} \frac{\sin x}{x(x+3)}\]We can write it like this,\[\Large\bf\sf =\lim_{x\to0}\frac{\sin x}{x}\cdot \lim_{x\to0}\frac{1}{x+3}\]

zepdrix (zepdrix):

Does that first term look familiar? It's an identity that you're supposed to remember! :D

OpenStudy (anonymous):

Yes. sinx/x equals one. Special trig identity. So the answer is 1/3?

zepdrix (zepdrix):

Yesssss, good job! :)

OpenStudy (anonymous):

Thank you so much.!!!